# A Non-perturbative General Approximation Scheme (NGAS) for Interacting Quantum Systems with Perturbation Theory for Arbitrary Strength of Interaction

### v2

@@ VIII.2 The simple harmonic-oscillator(SHO) in the ISW-approximation
The exact analytic results for the quantum-SHO are, well known and given in any standard text on quantum mechanics. There is therefore no need for any approximation for this celebrated example. However, these ‘exact’ results provide a further test of the accuracy and efficacy of the ISW-approximation in NGAS. It is purely for this reason that the SHO case is discussed here as a particular case ($g=0$) of the AHO.
The Hamiltonian of the SHO is simply obtained from that of the AHO, eqn.(51) by setting the quartic-coupling $g$ to zero and given by
@@
$n_{s}$ $E_{n}^{LO}$ $Error-LO{(\%)}$ $E_{n}^{(2)}$ $Exact$ $Error{(\%)}$
@@
In this context, it may be interesting to obtain an $asymptotic$ estimate for the accuracy of the LO-result given by eqn.(73) as compared to the `$exact^{\prime}$ result. This is given by
@@
which corresponds to an error of approximately $9.31\%$ At finite values of $n_{s}$ the errors are of the same order of magnitude.However, the inclusion of the 2nd-order correction in IPT significantly improves the accuracy of approximation.
@@ VIII.3 The QDWO
The Hamiltonian for the QDWO is given by (see, eqn.(36))
@@
Apart from the various applications of the DWO as discussed earlier, there is considerable theoretical/mathematical interest in the system. In particular, the instability at $g=0$ ( due to the non-existence of a physical ground state) prevents the application of the naïve perturbation theory. In some versions of modified-perturbation theory it has been established [95] that power series-expansion in $g$ is not even Borel-summable. However, in the context of NGAS, the ISW-approximation is routinely applicable to the DWO-case as well, by merely a change of sign of the $x^{2}$: $x^{2}\rightarrow~{}-x^{2}$ in the corresponding formulae for the AHO. In particular, the equation analogous to eqn.(63) for the DWO now becomes:
@@
leading to the $physical$ solution (analogous to eqn.(66)) given by
@@
[$~{}\left[\left\{\sqrt{\rho+1}+1\right\}^{\frac{1}{3}}-~{}\left\{\sqrt{\rho+1}-1\right\}^{\frac{1}{3}}~{}\right]~{}~{},$ (80)
@@
However, several authors [105] have found it convenient to measure the energy of the DWO from the bottom of the (symmetric) double-well, which means that a term equal to $1/{16g}$ be added to formula, eqn.(79). We denote this quantity as: $E_{n}^{\textrm{ref}}\equiv~{}E_{n}^{LO}+\left(1/{16g}\right)$.
In Table-III, we present the results for $E_{n}^{\textrm{ref}}$ along with results corrected to include 2nd-order perturbation effects in IPT ( see, Subsection-D below) for a range of values of $g$ and $n_{s}$. The relative accuracy of these results with respect to those from earlier computation are also given in this Table. Also shown, are corresponding results from ref.[105] denoted as $Exact$ under column-6.The entries in the last column correspond to (%$({\%})error$ of results under column-5 with respect to those under column-6. The entries under column-3 represent relative accuracy of the LO-results as compared with the $Exact$ results shown under column-6.
@@
[…]
It can be seen from this tabulation that the ISW-approximation uniformly reproduces the standard results to within $\sim{(2-10)~{}\%}$ for the case of LO and to within $\sim{(0.5-7.5)~{}\%}$ with the inclusion of 2nd-order correction in IPT, which is considered in details in the next subsection.
@@ VIII.4 Method for higher order corrections
In the context of NGAS, the ’improved perturbation theory (IPT)’ is the development of the standard Rayleigh-Schrödinger perturbation series (RSPS) with the choice of the $unperturbed~{}Hamiltonian$ as the input-Hamiltonian, $H_{0}$ with the free-parameters in the latter being determined through PEQA and CO (see,eqs.(16-17)). The $perturbation,H^{{}^{\prime}}$ is then defined by eqn.(21). The following properties of the IPT follow from the above assignment which are distinct from those in the conventional perturbation theory in the following aspects:
(i) Using eqn.(20) and eqn.(59) it is seen that the perturbation-correction remains always sub-dominant (by construction), compared to the unperturbed part in the following average-sense:
@@
This property may be contrasted with the corresponding situation in the $conventional$ RSPS. Consider e.g. the case of the AHO where the ‘perturbation’ ultimately prevails over the ’unperturbed’ component of the Hamiltonian, no matter however small the quartic-coupling is.
(ii) Secondly, the 1st-order perturbation-correction identically vanishes due to eqn.(20),
(iii) The IPT is not restricted to small values of the coupling strength $g$ since eqn.(80) holds for arbitrary values of the latter.
(iv) There is no small-parameter naturally associated with the perturbation-however, the latter remains small in the average sense as defined by eqn.(80). Therefore, to keep track of order-by-order corrections in IPT, one can adopt the standard trick of introducing an arbitrary, real but finite parameter, $\eta$ through the substitution: $H^{{}^{\prime}}\rightarrow~{}{\eta}H^{{}^{\prime}}$ and set this parameter to unity after the computation.
(v)Property (iv) further implies the ``$universality^{\prime\prime}$ of application of IPT to arbitrary interaction since the perturbation, $H^{{}^{\prime}}$ does not involve any parameter of the original Hamiltonian as the expansion-parameter for the RSPS.
Because of the property (ii) the first non-trivial correction for the $n^{th}$ energy level starts at the 2nd-order and given by the standard expression:
@@
In eqn.(81), we have used the notation:
[…]
where $H^{{}^{\prime}}$ is defined by eqn.(21) and we have displayed the $(n,g)$-dependence of the relevant quantities. As an example, the perturbation-part of the Hamiltonian for the case of the quartic-AHO is given by
@@
where $h(n,g)$ is given by eqn.(69).
It may be noted that the RSPS can be derived [106] without invoking the properties of the eigenfunctions of the unperturbed Hamiltonian. Therefore, the non-orthogonality properties as expressed in eqn.(70) do not affect the standard formule of the RSPS.
As has been discussed earlier and noted in the Tables-(I-III), the inclusion of the 2nd-order correction to the energy levels defined as: $E_{n}^{(2)}\equiv~{}E_{n}^{LO}~{}+~{}\Delta{E_{n}^{(2)}}$ significantly improves the accuracy of the approximation in all the cases considered here, viz the AHO, DWO and the SHO. (Computation of still higher-order corrections to energy-levels can be carried out by standard techniques in a straight-forward manner as discussed in Section-X).
@@ VIII.5 Section-VIII - Summary
In summary, we have presented a very simple yet accurate approximation in the framework of the NGAS for the quartic anharmonic and the double-well oscillators using the elementary system of the infinite-square-well(ISW) as the approximating Hamiltonian. In the computation of the energy-spectrum, uniform accuracy is achieved to within a few percent of the exact numerical results even in the leading-order, for arbitrary values of the quartic-coupling, $g$ and level index,$n_{s}$ for all the above systems. This situation may be contrasted with the results obtained using text-book methods of approximations, e.g. the naive perturbation method, the variational calculations, the WKB-method etc. Besides, the formalism naturally reproduces the correct analytic-structure of the energy in the complex-$g$ plane otherwise established through rigorous mathematical analysis. To systematically improve the leading-order results further, an improved perturbation theory is formulated in NGAS which is not restricted to small-coupling-expansion and shown to yield further significant improvement in accuracy with the inclusion of the 2nd-order correction only. The results and the method are directly relevant in a pedagogic-context owing to the extreme simplicity of the scheme (NGAS) as well as, the input ISW-approximation. The method can be extended in a straight forward manner to deal with the cases of higher anharmonicity.
@@ IX Application of the NGAS in the Leading Order to $g\phi^{4}$ QFT in the Massive Symmetric Phase
In view of the successful application of the NGAS to the AHO’s and DWO’s as described in earlier sections, it is but natural to extend the formalism to ${}g\phi^{4}~{}$ quantum field-theory and test the consequences of the scheme. It may be worth while pointing out here that the ${}g\phi^{4}~{}$- field theory in physical dimensions ( and in lower dimensions ) is an important physical system which finds crucial applications in diverse areas of physics, e.g., the standard model of particle-physics (Higgs-mechanism) , cosmology [87], condensed matter physics , phase-transitions and critical phenomena [70] etc. Besides, this theory provides the simplest theoretical laboratory for testing the various approximation schemes in quantum field-theory (QFT). Hence it becomes imperative to test the current approximation scheme NGAS by applying the same to $g\phi^{4}$- QFT.
We, therefore take up this task in the following subsections employing the harmonic approximation for the input/ approximating Hamiltonian (AH):
@@ IX.2 The approximating Hamiltonian
To formulate the NGAS for the above theory we choose an approximating Hamiltonian (AH) denoted by ${}{\cal H}_{0}~{}$ such that ideally the defining constraints as given by eqs.(16-17) are satisfied. However, unlike the case in quantum mechanics, the PEQA (eqn.(16)) is hard to implement in QFT as this would involve multi-particle-states in evaluating the quantum-averages (QA). Hence, on grounds of simplicity, we relax the condition, eqn.(16) by restricting the QA to be evaluated in the “ few-particle” states only. To be more specific, we proceed as follows.
As before, the “ Approximating Hamiltonian” (AH) ${}{\cal H}_{0}~{}$ is given by the following expression :
@@
In addition, ${}\xi_{t}\equiv\partial\xi/\partial t~{};~{}\xi_{\alpha}\equiv\partial\xi/\partial x^{\alpha}$ etc. Eqn.(86) is at once identified to be the Hamiltonian density of the hermitian scalar-field ${}\xi(\vec{x},t)~{}$. The ’free parameters’ introduced in the AH are: $M$, $\sigma$ and $h_{0}$, which are to be determined as per the prescribed procedure by applying eqn.(16-17). For that purpose, it is first necessary to obtain the spectrum of ${}H_{0}~{}$ which is done as follows.
The diagonalization of the AH given by eqn.(86) is straightforward by using the Fourier expansion in terms of creation- and annihilation operators :
@@
$[b(\vec{k}),b^{\dagger}(\vec{q})]~{}=~{}\Omega_{k}(M)\delta^{3}(\vec{k}-\vec{q}),$ (92)
@@
The spectrum of the states are analogously obtained and denoted by :${}|vac>,|\vec{p}>,~{}|\vec{p_{1}},~{}\vec{p_{2}}>,...$ etc where the effective vacuum state $|vac~{}>$ is defined by
@@
and the multi particle-states are generated by multiple application of the creation-operator ${}b^{\dagger}(\vec{p})~{}$ on ${}|~{}vac~{}>~{}$:
@@
The implementation of PEQA require the evaluation of the QA of monomials of the field ${}\phi(\vec{x},t)~{}$ such as : $<vac|~{}\phi^{n}(\vec{x},t)~{}|vac>~{}\equiv~{}<\phi^{n}(\vec{x},t)~{}>,$ $<\vec{p}|~{}\phi^{n}(\vec{x},t)~{}|\vec{p}>~{}$etc. We first turn to the evaluation of ${}<\phi^{n}(\vec{x},t)>~{}$. This is readily done using ${}{translational-invariance}~{}$ of the vacuum-state :${}|vac>~{}$ and the ETCR,  as given by eqn.(4.7). Some useful results thus obtained are given below :
$<\phi(\vec{x},t)>~{}=~{}\sigma~{},$ (98)
$<\phi^{2}(\vec{x},t)>~{}=~{}\sigma^{2}~{}+~{}\int\frac{d^{3}\vec{k}}{\Omega_{k}(M)}~{}\equiv~{}\sigma^{2}~{}+~{}I_{0}~{},$ (99)
$<\phi^{4}(\vec{x},t)>~{}=~{}\sigma^{4}~{}+~{}6\sigma^{2}I_{0}~{}+~{}3I_{0}^{2}~{},~{}etc.$ (100)
@@
$<\phi_{\alpha}^{2}(\vec{x},t)>~{}=~{}\int\frac{d^{3}\vec{k}}{\Omega_{k}(M)}|\vec{k}|^{2},$ (101)
$<\phi_{t}^{2}(\vec{x},t)>~{}=~{}\int\frac{d^{3}\vec{k}}{\Omega_{k}(M)}~{}\omega_{k}^{2}(M)$ (102)
@@
Eqn.(102) can be interpreted as the generation of the ‘mass-gap’(i.e. shift in the bare-mass) due to interaction. In analogy with the terminology used earlier, we refer eqs.(101-102) as the “gap-equation(GE)” of the theory. This equation plays crucial role in the subsequent discussions.
The physical consequences of the theory in the leading-order (LO), such as the spectrum, renormalization, stability properties and the structure of the effective vacuum are discussed in the following subsections.

### v3

@@ VIII.2 The simple harmonic-oscillator(SHO) in the ISW-approximation
The exact analytical solutions for the quantum Simple Harmonic Oscillator (SHO) are well-documented and can be found in any standard quantum mechanics textbook. Consequently, there is no necessity for approximation of this renowned example. Nevertheless, these ’exact’ results serve as a benchmark for evaluating the accuracy and effectiveness of the ISW approximation in NGAS. It is for this reason that the SHO is examined here as a specific instance ($g=0$) of the AHO. The Hamiltonian of the SHO is derived from that of the AHO, as presented in equation (51), by setting the quartic coupling $g$ to zero, and is expressed as:
@@
$n_{s}$ $E_{n}^{LO}$ $Error-LO{(\%)}$ $E_{n}^{(2)}$ ExactExact $Error{(\%)}$
@@
In this context, it may be interesting to obtain an asymptoticasymptotic estimate for the accuracy of the LO-result given by eqn.(73) as compared to the $exact^{\prime}$ result. This is given by
@@
which corresponds to an error of approximately $9.31\%$ . At finite values of $n_{s}$ the errors are of the same order of magnitude.However, the inclusion of the 2nd-order correction in IPT significantly improves the accuracy of approximation.
@@ VIII.3 The QDWO
The Hamiltonian for the QDWO is given by (refer to, eqn.(36))
@@
Apart from the various applications of the QDWO as discussed earlier, there is considerable theoretical/mathematical interest in the system. In particular, the instability at $g=0$ ( due to the non-existence of a physical ground state) prevents the application of the naïve perturbation theory. In some versions of modified-perturbation theory it has been established [95] that power series-expansion in $g$ is not even Borel-summable. However, in the context of NGAS, the ISW-approximation is routinely applicable to the DWO-case as well, by merely a change of sign of the $x^{2}$: $x^{2}\rightarrow~-x^{2}$ in the corresponding formulae for the AHO. In particular, the equation analogous to eqn.(63) for the QDWO now becomes:
@@
leading to the physicalphysical solution (analogous to eqn.(66)) given by
@@
$~\left[\left\{\sqrt{\rho+1}+1\right\}^{\frac{1}{3}}-~\left\{\sqrt{\rho+1}-1\right\}^{\frac{1}{3}}~\right]~~,$ (80)
@@
However, several authors [105] have found it convenient to measure the energy of the DWO from the bottom of the (symmetric) double-well, which means that a term equal to 1/16g1/{16g} be added to formula, eqn.(79). We denote this quantity as: $E_{n}^{\textrm{ref}}\equiv~E_{n}^{LO}+\left(1/{16g}\right)$.
In Table-III, we present the results for $E_{n}^{\textrm{ref}}$ along with results corrected to include 2nd-order perturbation effects in IPT ( see, Subsection-D below) for a range of values of $g$ and $n_{s}$. The relative accuracy of these results with respect to those from earlier computation are also given in this Table. Also shown, are corresponding results from ref.[105] denoted as $Exact$ under column-6.The entries in the last column correspond to $({\%})error$ of results under column-5 with respect to those under column-6. The entries under column-3 represent relative accuracy of the LO-results as compared with the $Exact$ results shown under column-6.
@@
Table 3: Results for the Energy-levels of the QDWO in ISW-approximation, for sample values of $n_{s}~$ andg~g~. For explanation of other entries in the Table, see text.
The tabulated data indicate that the ISW-approximation consistently replicates the standard results with a deviation of approximately $\sim{(2-10)~\%}$ for the leading order (LO) case. Furthermore, when incorporating the second-order correction in the improved perturbation theory (IPT), the deviation is reduced to approximately $\sim{(0.5-7.5)~\%}$. This aspect is examined in detail in the subsequent subsection.
@@ VIII.4 Method for higher order corrections
In the context of NGAS, the ’improved perturbation theory (IPT)’ represents an advancement of the standard Rayleigh-Schrödinger perturbation series (RSPS), wherein the unperturbedHamiltonianunperturbed~Hamiltonian is selected as the input-Hamiltonian, $H_{0}$, with its free parameters determined through PEQA and CO (refer to eqs. (16-17)). The $perturbation,H^{{}^{\prime}}$ is subsequently defined by eqn. (21). The following properties of the IPT emerge from this configuration, distinguishing it from conventional perturbation theory in several respects:(i) By employing equations (20) and (59), it becomes apparent that the perturbation correction remains consistently subdominant (by design) to the unperturbed component in the following average sense:
@@
This characteristic is in contrast to the analogous situation in the conventional Rayleigh-Schrödinger Perturbation Theory (RSPT). For example, consider the case of the Anharmonic Oscillator (AHO), where the ’perturbation’ eventually dominates the ’unperturbed’ component of the Hamiltonian, irrespective of the magnitude of the quartic coupling.   (ii) Secondly, the first-order perturbation correction identically vanishes due to equation (20).   (iii) TheiImproved perturbation theory (IPT) is not restricted to small values of the coupling strength $g$, as equation (80) is valid for arbitrary values of the latter.   (iv) There is no small parameter inherently associated with the perturbation; however, it remains small in the average sense as defined by equation (80). Therefore, to track order-by-order corrections in IPT, one can utilize the standard technique of introducing an arbitrary, real but finite parameter, $\eta$, through the substitution: $H^{{}^{\prime}}\rightarrow~{\eta}H^{{}^{\prime}}$, setting this parameter to unity post-computation.   (v) Property (iv) further implies the ”universality” of IPT’s application to arbitrary interactions, as the perturbation, $H^{{}^{\prime}}$, does not involve any parameter of the original Hamiltonian in contrast to the  expansion parameter for the RSPT.
Due to property (ii), the initial non-trivial correction for the $n^{th}$ energy level begins at the second order and is represented by the standard expression:
@@
In eqn.(81), the notation used is:
$\langle~n|H^{{}^{\prime}}|~m\rangle~\equiv~\int_{-a(n,g)}^{+a(n,g)}dx~{\phi_{n}^{\ast}(x)~H^{{}^{\prime}}(n,g)~\phi_{m}(x)}$~~ (84)
where $H^{{}^{\prime}}$ is defined by eqn.(21), and the $(n,g)$-dependence of the relevant quantities is displayed. As an example, the perturbation part of the Hamiltonian for the quartic-AHO case is given by
@@
where $h(n,g)$ is provided by eqn.(69).
It is noteworthy that the RSPT can be derived [106] without invoking the properties of the eigenfunctions of the unperturbed Hamiltonian. Consequently, the non-orthogonality properties as expressed in eqn.(70) do not impact the standard formulae of the RSPT.
As previously discussed and noted in Tables-(I-III), the inclusion of the second-order correction to the energy levels, defined as: $E_{n}^{(2)}\equiv~E_{n}^{LO}~+~\Delta{E_{n}^{(2)}}$, significantly enhances the accuracy of the approximation across all cases considered here, namely the AHO, DWO, and the SHO. (Computation of higher-order corrections to energy levels can be conducted using standard techniques in a straightforward manner as discussed in Section-X).
@@ VIII.5 Section-VIII - Summary
In summary, we have presented a simple yet accurate approximation within the framework of the NGAS for the quartic anharmonic and double-well oscillators, utilizing the elementary system of the infinite-square-well (ISW) as the approximating Hamiltonian. In the computation of the energy spectrum, uniform accuracy is achieved within a few percent of the exact numerical results even at the leading order, for arbitrary values of the quartic coupling, $g$, and level index, $n_{s}$, for all the aforementioned systems. This situation contrasts with the results obtained using textbook methods of approximation, such as the naive perturbation method, variational calculations, and the WKB method. Furthermore, the formalism naturally reproduces the correct analytic structure of the energy in the complex-$g$ plane, otherwise established through rigorous mathematical analysis. To systematically improve the leading-order results, an enhanced perturbation theory is formulated in NGAS, which is not restricted to small-coupling expansion and is shown to yield further significant improvement in accuracy with the inclusion of the second-order correction only. The results and the method are directly relevant in a pedagogic context due to the extreme simplicity of the scheme (NGAS) as well as the input ISW approximation. The method can be extended straightforwardly to address cases of higher anharmonicity.
@@ IX Application of the NGAS in the Leading Order to $g\phi^{4}$ QFT in the Massive Symmetric Phase
In light of the successful application of the NGAS to the AHO and DWO, as detailed in preceding sections, it is logical to extend this formalism to the $g\phi^{4}~$ quantum field theory and examine the implications of this approach. It is pertinent to note that the $g\phi^{4}~$ field theory, both in physical and lower dimensions, represents a significant physical system with critical applications across various domains of physics, such as the standard model of particle physics (Higgs mechanism), cosmology [87], condensed matter physics, phase transitions, and critical phenomena [70], among others. Furthermore, this theory serves as a fundamental theoretical framework for evaluating various approximation schemes in quantum field theory (QFT). Consequently, it is essential to assess the current approximation scheme, NGAS, by applying it to $g\phi^{4}$ QFT. We undertake this task in the subsequent subsections, utilizing the harmonic approximation for the input/approximating Hamiltonian (AH):
@@ IX.2 The approximating Hamiltonian
To develop the NGAS for the aforementioned theory, we select an approximating Hamiltonian (AH) denoted by ${\cal H}_{0}~$ such that, ideally, the defining constraints as specified by eqs.(16-17) are satisfied. However, unlike the scenario in quantum mechanics, the PEQA (eqn.(16)) is challenging to implement in QFT as it involves multi-particle states in evaluating the quantum averages (QA). Therefore, for simplicity, we relax the condition, eqn.(16), by restricting the QA to be evaluated in the ”few-particle” states only. Specifically, we proceed as follows. As previously, the ”Approximating Hamiltonian” (AH) ${\cal H}_{0}~$ is represented by the following expression:
@@
In addition, $\xi_{t}\equiv\partial\xi/\partial t~;~\xi_{\alpha}\equiv\partial\xi/\partial x^{\alpha}$ etc.
Eqn.(86) is immediately recognized as the Hamiltonian density of the hermitian scalar field $\xi(\vec{x},t)~$. The ’free parameters’ introduced in the AH are: $M$, $\sigma$, and $h_{0}$, which are to be determined according to the prescribed procedure by applying eqn.(16-17). For this purpose, it is first necessary to obtain the spectrum of $H_{0}~$, which is accomplished as follows. The diagonalization of the AH given by eqn.(86) is straightforward using the Fourier expansion in terms of creation and annihilation operators:
@@
$b(\vec{k}),b^{\dagger}(\vec{q})]~=~\Omega_{k}(M)\delta^{3}(\vec{k}-\vec{q}),$ (92)
@@
The spectrum of the states are analogously obtained and denoted by $|vac>,|\vec{p}>,~|\vec{p_{1}},~\vec{p_{2}}>,$... etc where the effective vacuum state |vac>|vac~> is defined by
@@
and the multi particle-states are generated by multiple application of the creation-operator $b^{\dagger}(\vec{p})~$ on |vac>~|~vac~>~:
@@
The implementation of PEQA require the evaluation of the QA of monomials of the field $\phi(\vec{x},t)~$ such as : $vac|~\phi^{n}(\vec{x},t)~|vac>~\equiv~<\phi^{n}(\vec{x},t)~>, <$$\vec{p}|~\phi^{n}(\vec{x},t)~|\vec{p}>~$etc. We first turn to the evaluation of $\phi^{n}(\vec{x},t)>~$. This is readily done using translationalinvariance~{translational-invariance}~ of the vacuum-state :$|vac>~$ and the ETCR,  as given by eqn.(4.7). Some useful results thus obtained are given below :
$\phi(\vec{x},t)>~=~\sigma~,$ (98)
$\phi^{2}(\vec{x},t)>~=~\sigma^{2}~+~\int\frac{d^{3}\vec{k}}{\Omega_{k}(M)}~\equiv~\sigma^{2}~+~I_{0}~,$ (99)
$\phi^{4}(\vec{x},t)>~=~\sigma^{4}~+~6\sigma^{2}I_{0}~+~3I_{0}^{2}~,~etc.$ (100)
@@
$\phi_{\alpha}^{2}(\vec{x},t)>~=~\int\frac{d^{3}\vec{k}}{\Omega_{k}(M)}|\vec{k}|^{2},$ (101)
$\phi_{t}^{2}(\vec{x},t)>~=~\int\frac{d^{3}\vec{k}}{\Omega_{k}(M)}~\omega_{k}^{2}(M)$ (102)
@@
Eqn. (102) can be interpreted as the generation of the ’mass-gap’ (i.e., shift in the bare mass) due to interaction. In analogy with the terminology used earlier, we refer to Eqs. (101-102) as the ”gap equation (GE)” of the theory. This equation plays a crucial role in the subsequent discussions. The physical implications of the theory in the leading order (LO), such as the spectrum, renormalization, stability properties, and the structure of the effective vacuum, are discussed in the following subsections.